OpenStudy (anonymous):

it IS EITHER C OR D PLEASE HELP Solve 2x2 − 8x = −7. negative 2 plus or minus square root of 2 negative 2 plus or minus 2 square root of 2 quantity of 2 plus or minus square root of 2 all over 2 2 plus or minus square root of 2 end root over 2

OpenStudy (anonymous):

HERE

OpenStudy (xapproachesinfinity):

complete the square, or use the quadratic formula!!

OpenStudy (xapproachesinfinity):

i would save time and do quadratic lolz

OpenStudy (welshfella):

try and solve it yourself - it will benefit you more than just knowing the answer

OpenStudy (anonymous):

It's D...am I right?

OpenStudy (xapproachesinfinity):

here is the quadratic if you don;t rem \(\Large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}~~~ for~a~general~equation~of ~the~form~ \\ax^2+bx+c=0\)

OpenStudy (xapproachesinfinity):

rewrite your equation like this \(2x^2-8x+7=0\) now you have everything under your nose just work :)